If in the same amount of time that it takes her to read 40 pages of economics and 30 pages of sociology, Martha could read 20 pages of economics and 110 pages of sociology, then which of these equations describes combinations of pages of economics, E, and sociology, S, that she could read in the time that it takes to read 40 pages of economics and 30 pages of sociology ?


A. E/2 = 50

B. 4E + S = 190

C. E + S = 70

D. E + S = 130

E. All of the above


1
Expert's answer
2021-09-27T09:07:30-0400

Option D is correct


Explanation: She could read 20 pages of E and 110 pages of S in the given time. So, E = 20 and S = 110.


So, E + S = 110+20 = 130